OpenStudy (anonymous):

m= - 1/4 and the coordinate (4,0)

OpenStudy (anonymous):

you want the equation?

OpenStudy (anonymous):

y= -1/4x +4???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the linear equation

OpenStudy (anonymous):

y=-1/4x+4

OpenStudy (anonymous):

howed you get that ?

OpenStudy (anonymous):

don't u substitute the points into y=ax+b?

OpenStudy (anonymous):

the formula y=mx+b plug in M (which you said was -1/4) and b which is (4,0)

OpenStudy (anonymous):

y=-1/4x-1, u need to substitute the points.

OpenStudy (anonymous):

(y1- y)=m(x1- x) plug in your number for y,x,m

OpenStudy (anonymous):

I use y=ax+b. much easier. just find b. and substitute y and x.

OpenStudy (anonymous):

it should be y=-x/4 +1

OpenStudy (anonymous):

oh i thought it just had to be put into point slope nvm XD

OpenStudy (anonymous):

THANKS YOU INIK!

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

see?! we were right! ahahahahahah!